# LongCaptioning: Unlocking the Power of Long Caption Generation in Large Multimodal Models

### v1

@@ 8.4 Frame-Level Captioning
Task: You are an expert in understanding the visual details of individual frames within a video. You are requested to create detailed descriptions for each video frame sent to you. Your task is to describe the frame’s content with high precision, focusing only on the elements visible in that exact frame. Do not infer or speculate about actions or events not explicitly visible in the frame. Guidelines For Frame Description: - Describe only what is visible in the frame: Focus on the exact visual details, without making assumptions about what happens before or after. - Avoid narrative progression: Unlike a clip description, there is no need to interpret or connect this frame with others. Only describe the current frame. - Be specific and exhaustive: Include as many details as possible, such as: - Objects: Colors, shapes, textures, positions, and relationships between objects. - People: Clothing, facial expressions, posture, gestures, and any visible features (e.g., hair color, accessories). - Background: Environmental details, lighting, shadows, and any visible text (with translations if necessary). - Text in the frame: If text appears, provide its original language and an English translation in parentheses. - No additional reasoning: Do not infer motivations, future actions, or unseen parts of the scene. Output Format: Your response should look like this: ”Frame Level Description”: ”The frame shows…”
8.5 Clip-Level Captioning
@@
Task: You are an expert in understanding scene transitions based on visual features in a video. There is a video including multiple sequential clips (clip-1,clip-2,…). Given the description for these clips (clip-1,clip-2,…,) as the context, you are requested to create the descriptions for the current clip sent to you, which includes multiple sequential frames. Guidelines For Clip Description: - Your description should see the description of previous clips as context. - Analyze the narrative progression implied by the sequence of frames, interpreting the sequence as a whole. - Note that since these frames are extracted from a clip, adjacent frames may show minimal differences. These should not be interpreted as special effects in the clip. - Note that some objects and scenes shown in the previous clips might not shown in the current clip. Be carefully do not assume the same object and scenes shown in every clips. - If text appears in the frames, you must describe the text in its original language and provide an English translation in parentheses. For example: book. Additionally, explain the meaning of the text within its context. - When referring to people, use their characteristics, such as clothing, to distinguish different people. - **IMPORTANT** Please provide as many details as possible in your description, including colors, shapes, and textures of objects, actions and characteristics of humans, as well as scenes and backgrounds. Output Format: Your response should look like this: ”Clip Level Description”: ”The clip begins with…, progresses by…, and concludes with…” Description of Previous Clips: {t-1_step_clip_description}
8.6 Video-Level Captioning
@@
Task: You are an expert at understanding frame-level and clip-level descriptions in a video that includes {num_frame} frames and {num_clip} clips. You are requested to create a video description by summarizing these frame-level and clip-level descriptions chronologically. Guidelines For Video Description: - Since the frame-level and clip-level descriptions are provided in chronological order, ensure that the video description is coherent and follows the same sequence. Avoid referring to the first or final frame of each clip as the first or final frame of the entire video. - Include any text that appears in the clip, provide its English translation in parentheses, and explain the significance of each text within its context. - The tone of the video description should be as if you are describing a video directly instead of summarizing the information from several clip descriptions. Therefore, avoid phrases found in the referred clip descriptions such as ”The clip begins…”, ”As the clip progresses…”, ”The clip concludes”, ”The final/first frame”, ”The second clip begins with”, ”The final frames of this segment”, etc - **IMPORTANT** Include all details from the given clip descriptions in the video description. Try to understand of the theme of the video and provide a coherent narrative that connects all the clips together. Output Format: 1. Your output should be formed in a JSON file. 2. Only provide the Python dictionary string. 3. You can use various descriptive sentence structures to outline the narrative progression. One example is: {} Your response should look like this: {{”Video Level Description”: ”YOUR DESCRIPTION HERE.”}} Frame-level Description (sorted in chronological order by number): {t_step_frame_descriptions} Clip-level Description (sorted in chronological order by number): {t_step_clip_description} Please give me the description of the video given the frame-level and clip-level descriptions.
Figure 7: Visualization of example.
Figure 8: Visualization of example.

### v2

@@ B.4 Frame-Level Captioning
Task: You are an expert in understanding the visual details of individual frames within a video. You are requested to create detailed descriptions for each video frame sent to you. Your task is to describe the frame’s content with high precision, focusing only on the elements visible in that exact frame. Do not infer or speculate about actions or events not explicitly visible in the frame. Guidelines For Frame Description: - Describe only what is visible in the frame: Focus on the exact visual details, without making assumptions about what happens before or after. - Avoid narrative progression: Unlike a clip description, there is no need to interpret or connect this frame with others. Only describe the current frame. - Be specific and exhaustive: Include as many details as possible, such as: - Objects: Colors, shapes, textures, positions, and relationships between objects. - People: Clothing, facial expressions, posture, gestures, and any visible features (e.g., hair color, accessories). - Background: Environmental details, lighting, shadows, and any visible text (with translations if necessary). - Text in the frame: If text appears, provide its original language and an English translation in parentheses. - No additional reasoning: Do not infer motivations, future actions, or unseen parts of the scene. Output Format: Your response should look like this: "Frame Level Description": "The frame shows…"
B.5 Clip-Level Captioning
@@
Task: You are an expert in understanding scene transitions based on visual features in a video. There is a video including multiple sequential clips (clip-1,clip-2,…). Given the description for these clips (clip-1,clip-2,…,) as the context, you are requested to create the descriptions for the current clip sent to you, which includes multiple sequential frames. Guidelines For Clip Description: - Your description should see the description of previous clips as context. - Analyze the narrative progression implied by the sequence of frames, interpreting the sequence as a whole. - Note that since these frames are extracted from a clip, adjacent frames may show minimal differences. These should not be interpreted as special effects in the clip. - Note that some objects and scenes shown in the previous clips might not shown in the current clip. Be carefully do not assume the same object and scenes shown in every clips. - If text appears in the frames, you must describe the text in its original language and provide an English translation in parentheses. For example: book. Additionally, explain the meaning of the text within its context. - When referring to people, use their characteristics, such as clothing, to distinguish different people. - **IMPORTANT** Please provide as many details as possible in your description, including colors, shapes, and textures of objects, actions and characteristics of humans, as well as scenes and backgrounds. Output Format: Your response should look like this: "Clip Level Description": "The clip begins with…, progresses by…, and concludes with…" Description of Previous Clips: {t-1_step_clip_description}
B.6 Video-Level Captioning
@@
Task: You are an expert at understanding frame-level and clip-level descriptions in a video that includes {num_frame} frames and {num_clip} clips. You are requested to create a video description by summarizing these frame-level and clip-level descriptions chronologically. Guidelines For Video Description: - Since the frame-level and clip-level descriptions are provided in chronological order, ensure that the video description is coherent and follows the same sequence. Avoid referring to the first or final frame of each clip as the first or final frame of the entire video. - Include any text that appears in the clip, provide its English translation in parentheses, and explain the significance of each text within its context. - The tone of the video description should be as if you are describing a video directly instead of summarizing the information from several clip descriptions. Therefore, avoid phrases found in the referred clip descriptions such as "The clip begins…", "As the clip progresses…", "The clip concludes", "The final/first frame", "The second clip begins with", "The final frames of this segment", etc - **IMPORTANT** Include all details from the given clip descriptions in the video description. Try to understand of the theme of the video and provide a coherent narrative that connects all the clips together. Output Format: 1. Your output should be formed in a JSON file. 2. Only provide the Python dictionary string. 3. You can use various descriptive sentence structures to outline the narrative progression. One example is: {} Your response should look like this: {{"Video Level Description": "YOUR DESCRIPTION HERE."}} Frame-level Description (sorted in chronological order by number): {t_step_frame_descriptions} Clip-level Description (sorted in chronological order by number): {t_step_clip_description} Please give me the description of the video given the frame-level and clip-level descriptions.
Figure 10: Qualitative comparison of different methods.
Figure 11: Qualitative comparison of LongCaptioning-8B and Ground Truth.